# DAOcc: 3D Object Detection Assisted Multi-Sensor Fusion for 3D Occupancy Prediction

### v2

@@ 4.4 Ablation Studies
pedestrian |
traffic cone |
✓ | 51.26 | 20.60 | 38.89 | 55.44 | 36.65 | |
✓ | ✓ | 52.82 (+1.56) | 36.35 (+15.75) | 45.48 (+6.59) | 58.27 (+2.83) | 41.98 (+5.33) |
44.31 | 14.94 | 32.22 | 43.77 | 34.07 | ||
✓ | 46.89 (+2.58) | 32.51 (+17.57) | 42.89 (+10.67) | 53.41 (+9.64) | 40.55 (+6.48) |
Ablation on the Auxiliary Detection Head. To fully leverage the inherent geometry and structure information within point cloud features, we incorporate 3D object detection as auxiliary supervision. As shown in Table 6, by adding auxiliary 3D object detection supervision, we can achieve an improvement of 1.56 mIoU when using camera visible masks during training and 2.58 mIoU when not using the mask, respectively. This demonstrates that adding a detection branch can significantly enhance the discriminability of the multi-modal features. Furthermore, we report on the four categories that benefit the most from the 3D object detection supervision. Notably, all four categories are foreground objects, which further validates the effectiveness and rationale behind the performance improvement achieved by our method. Specifically, when the camera visible mask is used during training, the addition of the auxiliary detection branch results in improvements of 15.75 mIoU, 6.59 mIoU, 2.83 mIoU, and 5.33 mIoU for bicycles, motorcycles, pedestrians, and traffic cones, respectively. Conversely, when the camera visible mask is not used during training, the improvements are 17.57 mIoU, 10.67 mIoU, 9.64 mIoU, and 6.48 mIoU for the same categories, respectively.
5 Conclusion
In this paper, we propose a novel multi-modal occupancy prediction framework, DAOcc, which aims to achieve superior performance while utilizing a deployment-friendly image backbone and practical input image resolutions. DAOcc achieves this superior performance by fully leveraging the unique advantages of point cloud data within a multi-modal framework, advantages that have been largely overlooked by previous methods. Extensive experiments conducted on the Occ3D-nuScenes and SurroundOcc datasets demonstrate the superiority of DAOcc compared to existing methods. We believe that DAOcc has practical application potential and could inspire future research in multi-modal occupancy prediction.
Supplementary Material
Method | Modality | GPU Type | FPS | mIoU |
FlashOcc(2f) [57] | C | A100-PCIE-40G | 0.9 | 43.52 |
RadOcc [60] | C+L | - | 0.3 | 49.38 |
EFFOcc [43] | C+L | RTX A6000 | 1.8 | 52.62 |
DAOcc(ours) | C+L | RTX 3070Ti | 2.7 | 53.82 |
6 Runtime Comparisons
We compare the runtime of different methods, as shown in Table 7. The FPS is obtained by testing on the corresponding GPU Type listed in the table. Since the source code for RadOcc is unavailable and EFFOcc has not yet released its implementation, we rely on the runtime speeds reported in their respective papers as benchmarks. It’s worth noting that the computing capability of the RTX 3070 Ti is significantly lower than that of both the RTX A6000 and the A100. The results demonstrate that our proposed DAOcc achieves the best performance while maintaining the fastest runtime speed.
7 Ablation for Achieving Final Performance
@@
In addition to the ablation studies presented in Section 4.4, this section provides further ablation experiments. Table 8 presents the ablation experiments that contributed to the final performance. First, by extending the training time from 15 epochs to 24 epochs, we can observe an improvement of 0.57 mIoU. Next, the use of CBGS [65] yields an additional 0.43 mIoU improvement. Since CBGS involves resampling, which increases the number of samples per epoch, we limit the training to 6 epochs to maintain a comparable total training time to the 24 epochs without CBGS for a fair comparison.
8 Detection Performance
@@
Since our proposed DAOcc incorporates an auxiliary 3D object detection branch into the occupancy prediction framework, we also evaluate the performance of this auxiliary branch. The auxiliary detection branch is designed to identify 10 object categories, including cars, trucks, buses, trailers, construction vehicles, pedestrians, motorcycles, bicycles, traffic cones, and barriers. All experiments are conducted using the nuScenes [3] dataset for the auxiliary 3D object detection training, and the results are evaluated using the mean average precision (mAP) and nuScenes detection score (NDS) metrics provided by the nuScenes benchmark.
As shown in Table 9, when DAOcc is trained on the Occ3D-nuScenes [44] dataset (for occupancy prediction) and the camera visible mask is used during training, the auxiliary detection branch achieves a performance of 59.4 mAP and 64.3 NDS. Notably, comparable performance is observed when the camera visible mask is not used. However, when the SurroundOcc [51] dataset is employed for occupancy prediction training, the auxiliary detection branch achieves a lower performance of 53.6 mAP and 60.2 NDS. We attribute this performance gap to the fact that the surroundocc dataset is automatically generated without manual calibration, potentially introducing more noise and negatively affecting model convergence.
EP15 | EP24 | CBGS (EP6) | mIoU |
✓ | 52.82 | ||
✓ | 53.39 | ||
✓ | 53.82 |
9 Visualization
@@
Dataset | Camera Mask | mIoU | mAP | NDS | Car | Truck | Bus | T.L. | C.V. | Ped. | M.C. | B.C. | T.C. | B.R. |
Occ3D-nuScenes [44] | ✓ | 53.8 | 59.4 | 64.3 | 85.7 | 57.2 | 65.8 | 33.1 | 22.8 | 75.5 | 68.3 | 47.3 | 70.2 | 68.2 |
Occ3D-nuScenes [44] | ✗ | 47.9 | 59.6 | 64.2 | 85.4 | 55.0 | 64.6 | 34.1 | 23.1 | 76.3 | 67.0 | 50.7 | 70.8 | 69.0 |
SurroundOcc [51] | - | 30.5 | 53.6 | 60.2 | 81.9 | 42.7 | 56.0 | 26.3 | 19.7 | 72.6 | 57.7 | 43.8 | 66.7 | 68.1 |

### v3

@@ IV Experiments
building |
vegetation |
tree trunk |
road |
sidewalk |
| BEVDet [16] | 9.88 | C | - | - | 0.13 | 13.06 | 2.17 | 10.15 | 7.80 | 5.85 | 4.62 | 0.94 | 1.49 | 0.0 | 7.27 | 10.06 | 2.35 | 48.15 | 34.12 |
| TPVFormer [40] | 16.76 | C | - | - | 3.89 | 17.86 | 12.03 | 5.67 | 13.64 | 8.49 | 8.90 | 9.95 | 14.79 | 0.32 | 13.82 | 11.44 | 5.8 | 73.3 | 51.49 |
| BEVFormer [39] | 16.76 | C | - | - | 3.48 | 17.18 | 13.87 | 5.9 | 13.84 | 2.7 | 9.82 | 12.2 | 13.99 | 0.0 | 13.38 | 11.66 | 6.73 | 74.97 | 51.61 |
| CTF-Occ [15] | 18.73 | C | 640 960 | R101-DCN | 6.26 | 28.09 | 14.66 | 8.22 | 15.44 | 10.53 | 11.78 | 13.62 | 16.45 | 0.65 | 18.63 | 17.3 | 8.29 | 67.99 | 42.98 |
| OPUS [65] | 19.00 | C | - | - | 4.66 | 27.07 | 19.39 | 6.53 | 18.66 | 6.41 | 11.44 | 10.40 | 12.90 | 0.00 | 18.73 | 18.11 | 7.46 | 72.86 | 50.31 |
| ODG [66] | 21.35 | C | 640 960 | R50 | 5.09 | 31.34 | 22.4 | 19.06 | 15.24 | 6.09 | 12.51 | 12.77 | 13.59 | 0.00 | 21.49 | 17.89 | 8.37 | 78.19 | 56.28 |
| LiDAR-Only [15] | 29.74 | L | - | - | 1.01 | 57.41 | 35.31 | 20.33 | 11.7 | 13.01 | 36.21 | 7.81 | 0.13 | 0.0 | 57.83 | 54.71 | 27.07 | 69.15 | 54.47 |
| BEVFormer-Fusion [15] | 39.05 | C+L | - | - | 5.11 | 64.61 | 52.35 | 21.52 | 32.74 | 17.1 | 42.62 | 27.75 | 13.36 | 0.05 | 63.65 | 60.51 | 35.64 | 81.89 | 66.84 |
| DAOcc (Ours) | 44.69 | C+L | 256 704 | R50 | 7.64 | 65.60 | 54.05 | 43.71 | 52.78 | 32.05 | 46.96 | 32.68 | 31.68 | 0.68 | 55.94 | 60.19 | 42.59 | 79.62 | 64.09 |
| DAOcc* (Ours) | 45.13 | C+L | 256 704 | R50 | 8.15 | 65.90 | 54.40 | 43.36 | 52.86 | 32.93 | 47.54 | 33.54 | 32.29 | 0.67 | 57.49 | 60.56 | 42.86 | 79.78 | 64.69 |
IV-A Datasets and Evaluation Metrics
IV-A1 Occ3D-nuScenes
Occ3D-nuScenes [15] is a large-scale benchmark dataset for 3D occupancy prediction, which contains 700 scenes for training and 150 scenes for validation. Each scene lasts for 20 seconds, with annotations provided at a frequency of 2 Hz. The dataset covers a perception range of [-40m, 40m] in the and directions, and [-1m, 5.4m] along the -axis, all within the ego coordinate system. This range is discretized into voxels, each with a size of [0.4m, 0.4m, 0.4m]. Each occupied voxel is assigned one of 17 semantic classes, including 16 common classes and a general object class labeled as ‘others’. The dataset includes RGB images captured by six surrounding cameras and LiDAR point clouds, enabling dense voxel-wise annotations for 3D scene understanding. Additionally, the dataset provides visibility masks for both LiDAR and camera modes, which can be utilized for training purposes.
IV-A2 Occ3D-Waymo
Occ3D-Waymo [15] is a large-scale benchmark dataset for 3D occupancy prediction, containing 798 training scenes and 202 validation scenes. It shares the same perception range and voxel size as Occ3D-nuScenes. Each occupied voxel is annotated with one of the 15 semantic classes, including a General Object (GO) class. The dataset provides RGB images from five cameras (with raw resolutions of for front, front-left, and front-right cameras, and for side-left and side-right cameras) and LiDAR point clouds, along with visibility masks for LiDAR and camera modalities to facilitate training.
IV-A3 Evaluation metrics
Following previous methods, we use the mean Intersection over Union (mIoU) as our evaluation metric when training models with the camera visible mask provided by Occ3D-nuScenes and Occ3D-Waymo, ensuring a fair comparison with existing approaches. When the camera visible mask is excluded from the training phase, we report an additional semantic segmentation metric, RayIoU [42], which is a recently introduced metric. RayIoU evaluates occupancy along query rays at three distance thresholds (1m, 2m, 4m), with the mean of these results serving as the final score.
IV-B Implementation Details
We use ResNet-50 [21], pre-trained on nuImages [24] by MMDetection [67], as our image backbone and adjust the input image resolution to 256×704. In the LiDAR branch, we aggregate LiDAR sweeps from 10 adjacent timestamps as input and initiate training from scratch. Data augmentation techniques, such as scaling, rotation, and flipping, are applied to the input images, and we also perform random flipping in the BEV space. The loss weights are configured as and . We adopt the AdamW optimizer [68] with a cosine annealing learning rate scheduler, incorporating a warmup phase, and set the initial learning rate to 2e-4. Unless conducting ablation experiments, we train our DAOcc using CBGS [69] for 6 epochs, the input point cloud range is set to [-54.0, -54.0, -5.0, 54.0, 54.0, 3.0] and the voxel size is set to [0.075, 0.075, 0.2]. All models are trained on 4 NVIDIA RTX4090 GPUs with a batch size of 4. Additionally, if the camera visible mask is not used during training, we randomly discard 80% of empty voxels, consistent with Zhao et al. [70]. Temporal information fusion and test-time augmentation are not used in any experiment. Unless otherwise stated, exponential moving average (EMA) for weight parameter is not used.
IV-C Comparison with State-of-the-Art Methods
IV-C1 Comparison on Occ3D-nuScenes with camera mask
As presented in Table I, we compare our proposed DAOcc with both image-based and multi-modal methods on the Occ3D-nuScenes dataset. All results are either provided directly by their authors or obtained from their official code implementations. FPS for SDGOCC and ALOcc are sourced directly from their original publications, evaluated on NVIDIA RTX 4090. All other FPS results are obtained through our own benchmarking on identical hardware (NVIDIA RTX 4090) using the PyTorch FP32 backend. Our proposed DAOcc establishes a new state-of-the-art performance, achieving an impressive 54.33 mIoU with only a 2D ResNet50 [21] backbone and a low input image resolution of 256×704, while maintaining a competitive inference speed of 7.8 FPS. Compared with the latest state-of-the-art multi-modal method SDGOCC, DAOcc achieves an improvement of 2.67 mIoU and maintains a higher FPS (7.8 vs 7.5). Furthermore, DAOcc performs strongly across nearly all foreground semantic categories, which are critical for ensuring autonomous driving safety.
As shown in Table II, we further evaluate the performance and inference speed of DAOcc under different frameworks and hardware configurations. Unlike most LSS-based methods that rely on custom CUDA operators (e.g., BEV Pool [49]), the image branch of DAOcc does not require any custom operators when deployed with TensorRT and can be directly exported to the ONNX format, greatly simplifying the deployment process. With TensorRT optimization, DAOcc achieves 104.9 FPS on an RTX 4090 while maintaining 54.2 mIoU, and 20.0 FPS on a Jetson AGX Orin (64GB) while maintaining 53.7 mIoU. Although quantization introduces a slight drop in mIoU performance, the resulting 53.7 mIoU still significantly surpasses all other methods reported in Table I.
IV-C2 Comparison on Occ3D-nuScenes without camera mask
To the best of our knowledge, no existing multi-modal occupancy prediction methods have reported their performance using the RayIoU metric, so we mainly compare with image-based methods. It is worth noting that some image-based methods [16, 17, 62, 44, 8], such as the current state-of-the-art STCOcc and ALOcc, are based on Explicit Depth Supervision [38] and leverage depth information derived from point clouds to supervise monocular depth estimation during training. As such, these methods can also be considered multi-modal in the training phase.
As shown in Table III, benefiting from the effective use of multi-modal inputs, especially point cloud data, our DAOcc achieves 48.3 mIoU and 48.4 RayIoU, significantly outperforming image-based methods by a large margin (surpassing ALOcc by 10.3 mIoU and 4.7 RayIoU), despite many of these methods employing multi-frame temporal fusion, which is an effective technique to improve RayIoU scores (e.g., +3.3 RayIoU for Panoptic-FlashOcc with 8-frame fusion vs. single-frame). Furthermore, when using the same 2D backbone (ResNet50) and image input resolution (256×704), although DAOcc additionally inputs point cloud data, its inference speed is still faster than the image-based method ALOcc (7.8 vs. 6.0).
IV-D Ablation Studies
Ablation studies are conducted on the validation set of Occ3D-nuScenes and Occ3D-Waymo, with the camera visible mask being used during training. Unless otherwise specified, all experiments presented in this section are trained for 15 epochs without employing CBGS [69].
IV-D1 Effectiveness of the BVRE
We propose the BVRE strategy to enrich spatial contextual information by expanding the perceptual range from a BEV (Bird’s Eye View) perspective. As shown in Table IV, a comparison of (c) and (d) indicates that, with the same voxel size of and the H-Range of [-5m, 3m], expanding the BEV range from [-41.4m, 41.4m] to [-45.6m, 45.6m] yields a performance improvement of 0.22 mIoU (from 50.76 to 50.98). Furthermore, the comparison of (c) and (g) shows that further expanding the BEV range to [-54.0m, 54.0m] with the same voxel size and H-Range increases the performance gain to 0.5 mIoU (from 50.76 to 51.26). This observation validates the value of spatial context information in a larger BEV range.
In addition, by comparing (d) and (e), the voxel size has a great impact on the model’s performance. Specifically, increasing the voxel size from to while keeping the BEV range and H-Range the same leads to a decrease in mIoU from 50.98 to 49.75. Moreover, by analyzing the data in (a), (b), and (c), the H-Range of [-5m, 3m] is a relatively optimal setting, as it consistently achieves higher mIoU values compared to other H-Ranges. To achieve better performance, we adopt the setting (g) in DAOcc, which combines the largest BEV range of [-54.0m, 54.0m], the voxel size of m, and the H-Range of [-5m, 3m].
IV-D2 Ablation on the Auxiliary Detection Head
To fully leverage the inherent geometry and structure information in point cloud features, we incorporate 3D object detection as auxiliary supervision. As shown in Table V, adding auxiliary 3D detection supervision leads to a 1.56 mIoU improvement when training with camera visible masks. Notably, this improvement originates entirely from foreground categories, which are highlighted in light gray in Table V. These categories correspond to the ones supervised by the auxiliary detection branch, demonstrating the effectiveness of our proposed supervision strategy.
Meanwhile, we observe that categories without detection supervision all exhibit varying degrees of performance degradation. We infer that this is because detection supervision emphasizes the localization of foreground objects, which strengthens the representation of foreground objects and thus weakens that of background categories. To mitigate this issue, we believe that additional BEV segmentation supervision could be introduced to balance the representation of foreground and background categories.
To further investigate how auxiliary 3D detection supervision affects the BEV feature distribution, we visualize intermediate features in Figure 5. Subfigures (a) and (b) show the input point cloud and the corresponding ground-truth occupancy map, while (c) and (d) present the BEV features without and with auxiliary 3D detection supervision, respectively. The comparison reveals that incorporating auxiliary 3D detection enhances the preservation of structural scene details and significantly improves the distinguishability of foreground objects.
IV-D3 Ablation on Loss Weight and LiDAR Sweeps
Figure 6 illustrates the impacts of two factors: the loss weight and LiDAR sweeps on the mIoU.
To analyze the effect of loss weight , which balances the loss contributions between 3D occupancy prediction and 3D object detection, we vary its value across 0.1, 0.01, 0.001, and 0.0001. As shown in the left plot, mIoU rises sharply from 48.17 to 52.82 as decreases from 0.1 to 0.01, then gradually declines to 50.81 as is further reduced to 0.0001. This suggests that there exists an optimal (around 0.01) that best balances the two loss terms for optimal mIoU.
The right plot explores the effect of LiDAR sweeps, referring to the number of point cloud frames aggregated over time. With sweeps increasing from 1 to 9, mIoU climbs steadily: starting at 50.91 (1 sweep), reaching 52.09 (3 sweeps), maintaining stability up to 5 sweeps, and then rising to 52.82 at 9 sweeps. Incorporating more sweeps generally enriches features from multi-frame data and improves mIoU performance, although the marginal gains diminish beyond a certain point.
IV-D4 Ablation for Achieving Final Performance
@@
Table VI presents the ablation experiments that contributed to the final performance of DAOcc. First, extending the training time from 15 epochs to 24 epochs yields an improvement of 0.57 mIoU. Next, using CBGS [69] yields an additional 0.43 mIoU improvement. Since CBGS employs resampling, which increases the number of samples per epoch, we limit the training to 6 epochs to maintain a total training time comparable to that of the 24 epochs without CBGS for a fair comparison. Finally, increasing the learning rate can further improve performance.
IV-D5 Comparison on Occ3D-Waymo with camera mask
We further evaluate the robustness and generalization capability of DAOcc on the Occ3D-Waymo dataset, which remains a relatively underexplored benchmark for occupancy prediction. Following prior works [15, 65, 66], we train DAOcc using only 20% of the training data and discard LiDAR sweeps by utilizing a single LiDAR frame for each input, ensuring a fair comparison. As presented in Table VII, DAOcc significantly outperforms all existing methods, surpassing the multi-modal baseline BEVFormer-Fusion by 5.64 mIoU. Compared to the nuScenes dataset, which features a single top-mounted LiDAR, the Waymo dataset provides denser point clouds captured from five LiDARs. This richer sensory input further amplifies the performance gap between multi-modal and image-only methods. For instance, the recently proposed image-based method ODG [66] achieves an mIoU that is only 47.8% of that achieved by DAOcc (21.35 vs. 44.69). These results highlight the critical importance of high-quality point cloud data in improving 3D occupancy prediction, particularly in complex environments, while also validating the strong generalization ability of DAOcc across different datasets and sensing setups.
IV-D6 Robustness
We aim to apply state-of-the-art 3D occupancy prediction models to real-world scenarios by leveraging sparse LiDAR data more effectively within a multi-modal fusion framework. Although robust alignment and fusion of sensor modalities are not the primary focus of this work, we nonetheless evaluate DAOcc’s robustness under challenging scenarios, as shown in Figure 7.
In the nighttime scenario, under low illumination, the vehicle ahead (highlighted by the yellow dashed box) is barely observable in the camera image except for its lights. Even so, DAOcc successfully recognizes this vehicle and reconstructs a voxelized occupancy shape that closely matches the ground truth, demonstrating strong robustness in low-light environments.
In the rainy scenario, raindrops obscure the camera view. The vehicle parked on the left (yellow dashed box) is nearly invisible, yet DAOcc still recognizes it accurately. A pedestrian partially occluded by another vehicle on the right (yellow dashed circle) is also correctly reconstructed in voxel space. Although DAOcc misses the vehicle immediately behind this pedestrian, this target is almost indiscernible even by manual inspection, and the corresponding LiDAR data also does not show vehicle traces at this position. We thus attribute this miss to the intrinsic difficulty of the sample rather than a failure of our method.
The above results indicate that although DAOcc was not specifically designed for robustness, it already exhibits inherent robustness against interference due to the intrinsic generalization of data-driven learning paradigms. Based on this observation, we believe that increasing training data and designing specific data augmentation strategies for simulating sensor failures during training would further enhance DAOcc’s robustness.
IV-D7 Detection Performance
@@
Since our proposed DAOcc incorporates an auxiliary 3D detection branch into the occupancy prediction framework, we also evaluate the performance of this auxiliary branch. The auxiliary detection branch identifies 10 object categories, including cars, trucks, buses, trailers, construction vehicles, pedestrians, motorcycles, bicycles, traffic cones, and barriers. We train the auxiliary 3D detection branch on the nuScenes dataset [24], and the results are evaluated using the mean average precision (mAP) and nuScenes detection score (NDS) metrics provided by the nuScenes benchmark. DAOcc achieves 59.4 mAP and 64.3 NDS on the nuScenes validation set, and the APs for the 10 object categories (in the order listed above) are 85.8, 55.5, 66.1, 30.8, 22.0, 75.8, 67.4, 52.2, 69.6, 69.2. These results show that DAOcc has the potential to be extended into a multi-task 3D perception framework.
IV-E Visualization
@@
V Conclusion
In this paper, we propose a novel multi-modal occupancy prediction framework, DAOcc, which aims to achieve superior performance while maintaining a deployment-friendly architecture with practical image backbones and input image resolutions. DAOcc achieves superior performance by fully leveraging the inherent geometry advantages of LiDAR data within a multi-modal framework, advantages largely overlooked by previous work. Extensive experiments on the Occ3D-nuScenes and Occ3D-Waymo datasets demonstrate that DAOcc surpasses existing methods in both accuracy and inference speed. Furthermore, due to its deployment-friendly design, DAOcc can be efficiently deployed and evaluated on edge devices. Notably, DAOcc runs in real-time on an AGX Orin device while maintaining state-of-the-art mIoU performance.